Using the equation for the Inflation Expectation Augmented Phillips Curve calculate and explain the expected level of wage inflation under the following three scenarios: µ = 6%, µ = 3%, µ = 5% LET: µ*= 5%, ε = .75, πe =4%. In your answer consider expected inflation and the unemployment gap


1
Expert's answer
2021-04-09T06:57:46-0400

Wage level inflation

Use formula m=πe -ε (u*-u)

M= 4-0.75(5-4.7)=3.775%

Unemployment gap

3.775-0.75=3.025%